# Fairness-aware Contextual Dynamic Pricing with Strategic Buyers

### v1

@@ Abstract
Contextual pricing strategies are prevalent in online retailing, where the seller adjusts prices based on products’ attributes and buyers’ characteristics. Although such strategies can enhance seller’s profits, they raise concerns about fairness when significant price disparities emerge among specific groups, such as gender or race. These disparities can lead to adverse perceptions of fairness among buyers and may even violate the law and regulation. In contrast, price differences can incentivize disadvantaged buyers to strategically manipulate their group identity to obtain a lower price. In this paper, we investigate contextual dynamic pricing with fairness constraints, taking into account buyers’ strategic behaviors when their group status is private and unobservable from the seller. We propose a dynamic pricing policy that simultaneously achieves price fairness and discourages strategic behaviors. Our policy achieves an upper bound of $O(\sqrt{T}+H(T))$ regret over $T$ time horizons, where the term $H(T)$ arises from buyers’ assessment of the fairness of the pricing policy based on their learned price difference. When buyers are able to learn the fairness of the price policy, this upper bound reduces to $O(\sqrt{T})$. We also prove an $\Omega(\sqrt{T})$ regret lower bound of any pricing policy under our problem setting. We support our findings with extensive experimental evidence, showcasing our policy’s effectiveness. In our real data analysis, we observe the existence of price discrimination against race in the loan application even after accounting for other contextual information. Our proposed pricing policy demonstrates a significant improvement, achieving 35.06% reduction in regret compared to the benchmark policy.
@@ 1 Introduction
Practical scenarios often involve price discrimination against specific groups, even after controlling for contextual information (Bocian et al., 2008; Zhang, 2018; Bartlett et al., 2022; Butler et al., 2023). One such example is the mortgage market. Our motivation comes from the Home Mortgage Disclosure Act (HMDA) data [^4] https://ffiec.cfpb.gov/data-publication/2022, where Popick (2022) found persistent group-level pricing disparities between minority applicants and borrowers from the majority race, even after accounting for credit risk factors. This suggests that minority applicants pay higher interest rates, even when other credit risk factors are held constant, underscoring the unfairness of such practices. We refer to Section 6 for more discussion of this HMDA dataset. Another instance can be seen in auto repair industry, where discrimination against female customers is prevalent (Busse et al., 2017). Nationwide, women are commonly charged more than men for the same auto repair work. In Los Angeles, 20% of auto shops surveyed quoted higher prices for women. On average, women are charged 8% more than men for repair jobs across the country [^5] https://abc7.com/women-overcharged-in-auto-repair-shops-charges-charged-for-repairs-pal/1660671/.
Unfairness in the pricing policy not only can lead to losses for the seller, but also can provoke strategic behaviors among buyers. In personalized pricing, buyers should not be able to easily obtain prices intended for a different consumer group, or if they can, the process should be sufficiently costly (Lukacs et al., 2016). Our study delves into buyers’ strategic behaviors, where such behavior is defined as buyers pretending to belong to an alternative group, incurring a fixed cost in the process (Li and Li, 2023). As revealed in our analysis of HMDA data in Section 6, minority applicants tend to pay higher interest rates compared to the majority group. In response, applicants from the minority group may strategically manipulate their identity to appear as members of the majority group, aiming for a lower interest rate. Practical instances of such strategic behaviors exist in reality. For instance, homeowners from the minority race asked friends from the majority race to pretend to be homeowners during property appraisals, leading to a significant increase in property value [^6] https://www.cnn.com/2021/12/09/business/black-homeowners-appraisal-discrimination-lawsuit/index.html
@@
. In cases like these, sellers often obtain buyers’ race information on the basis of buyers’ provided information, visual observation or surname [^7] https://www.consumerfinance.gov/rules-policy/regulations/1003/b/. Disadvantaged buyers may seek the assistance of more advantaged friends to appear in the process of purchasing. Other strategic behaviors include manipulating device information in the presence of price discrimination on device information such as Orbitz’ s price discrimination on Mac users (Mattila and Choi, 2014), forging a student ID if there is a discount for students. When significant unfairness arises, buyers may be motivated to manipulate their group membership to gain access to a lower price, even if it incurs additional costs.
@@ 1.1 Our Contribution
To address the aforementioned contextual dynamic pricing problem with fair-minded and strategic buyers, in this paper we propose a fairness-aware pricing policy designed to deter buyers’ strategic actions by fostering a favorable fairness perception.
@@ 1.2 Literature Review
Dynamic Pricing with Fairness. Dynamic pricing has been an active research area in operations research and machine learning (Luo et al., 2022; Fan et al., 2024). In the online pricing realm, Xu et al. (2023); Chen et al. (2023b); Cohen et al. (2024) have explored the non-contextual pricing problem with fairness constraint. Specifically, Chen et al. (2023b) and Cohen et al. (2024) examined pricing problems where the identical deterministic price is offered within each buyer group, while Xu et al. (2023) introduced random prices generated by probability distributions within each buyer group. Consequently, in these studies, buyers from the same group are subject to the same price or prices from the same probability distribution. In contrast, our work integrates contextual information into the pricing policy, offering prices based on features while simultaneously ensuring fairness among buyers with the same features from different groups. The work by Chen et al. (2023a) studied the contextual pricing problem with fairness constraint. In all the aforementioned literature, the true group status is observable by the seller while the true group status is unobservable in our paper. Moreover, we consider inequity-averse buyers who actively seek lower prices by manipulating their group status - an aspect not addressed in Xu et al. (2023); Chen et al. (2023a, b); Cohen et al. (2024). Another critical difference is that the aspect of how buyers learn about price fairness is overlooked in Xu et al. (2023); Chen et al. (2023a, b); Cohen et al. (2024). In contrast, our approach delves into the intricacies of how buyers learn about fairness. The existence of buyers’ strategic behaviors is well motivated in many pricing applications, rendering existing pricing tools not applicable. To address this challenge, we must devise new tools capable of handling both strategic behaviors and the learning process of fairness.
Dynamic Pricing with Strategic Buyers. Existing literature on pricing with strategic buyers has primarily focused on timing (Chen and Farias, 2018), untruthful bidding in pricing and auction design (Amin et al., 2014; Mohri and Munoz, 2015), and feature manipulation (Liu et al., 2024). Other literature also exists on feature manipulation within classification (Hardt et al., 2016; Dong et al., 2018; Chen et al., 2020; Ghalme et al., 2021; Bechavod et al., 2021). Our work specifically addresses strategic behaviors related to feature manipulation, closely related to the study by Liu et al. (2024). However, the policy presented by Liu et al. (2024) incurs a social loss as it cannot effectively curb the futile strategic manipulation. In contrast, our paper employs fairness as a tool to discourage strategic behaviors. Furthermore, Liu et al. (2024) could not enforce fairness in pricing policies and hence is not applicable to address price discrimination.
@@ 1.3 Organization
The remainder of the paper is organized as follows. In Section 2, we introduce the new problem and the necessary components: fairness and strategic behaviors. In Section 3, we propose the fairness-aware pricing policy with strategic buyers. In Section 4, we provide the theoretical analysis. In Section 5, we conduct simulation studies to examine our proposed policy. A real data analysis is provided in Section 6, followed by some discussion of future work in Section 7. All proofs are included in the supplementary materials.
@@ 2 Problem Formulation
We first introduce the setting of the contextual dynamic pricing problem. At each time $t$, a buyer with feature $\bm{x}_{t}\in\mathbb{R}^{d}$ and a group status $G_{t}\in\{0,1\}$ enters the market. The true group status of each buyer is considered as a private type, which is unobservable by the seller. The buyer can manipulate it to the other group by incurring some cost. We denote $C_{0}>0$ as the fixed minimum unit cost of manipulation, averaged over the total number of products purchased. Upon receiving $\bm{x}_{t}$ and a reported group status $G_{t}^{\prime}$, which may be different from $G_{t}$, the seller offers a price $p_{t}$. At time $t$, the demand of a buyer with feature $\bm{x}_{t}$ and group status $G_{t}=j\in\{0,1\}$ is
$y_{jt}=\alpha_{j}p_{jt}+\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t}+\epsilon_{t},$ (1)
[…]
Now we consider the dynamic pricing problem with price fairness. To incorporate the price fairness, we consider a price constraint, $|p_{0t}-p_{1t}|\leq\delta$, where $\delta\geq 0$ is the parameter for the fairness level that is selected by the seller to meet the internal goal of the company or satisfy regulatory requirements. This price constraint indicates that the price difference of the buyers from two different groups should not exceed $\delta$ after controlling other features.
@@
The objective function of the weighted revenue in (2) is an extension of Xu et al. (2023) from the non-contextual setting to the contextual setting. It also includes Cohen et al. (2024); Chen et al. (2023b) as a special case with $q=1/2$.
The constrained optimization problem (2) serves as a full-information benchmark for evaluating our pricing policy. By solving (2) as detailed in Appendix S.1, we obtain the optimal prices $p_{0t}^{*}=p_{0}^{*}(\bm{x}_{t})$ and $p_{1t}^{*}=p_{1}^{*}(\bm{x}_{t})$, i.e., for $j=0,1$,
@@
Now, we discuss buyers’ strategic behavior. The group status of the buyer is considered as a private type and is unobserved. The buyers can strategically deceive the seller about the group status to pursue a lower price. Remind that the cost of manipulating group status is $C_{0}>0$, which is public information. Given that group 0 is the discriminated group, only the buyers from group 0 are likely to manipulate the group status. Let $\widehat{p}_{0}(\bm{x})$ and $\widehat{p}_{1}(\bm{x})$ be the prices for group 0 and group 1 that the buyer has estimated using the history data. For a buyer from group 0 with feature $\bm{x}$, the total estimated cost is $\widehat{p}_{0}(\bm{x})$ without group manipulation and $\widehat{p}_{1}(\bm{x})+C_{0}$ after group manipulation. Therefore, the buyer from group 0 will strategically report the group status as
@@ 2.1 Linear Regret for Existing Fair Pricing Policy
Consideration of buyers’ fairness learning is critical in the dynamic pricing with strategic buyers. Existing fair pricing policies (Chen et al., 2023a, b; Cohen et al., 2024) do not consider buyers’ strategic behaviors and ignore buyers’ fairness learning process. In this case, even the seller provides prices with fairness constraints, the disadvantaged buyers always act strategically by manipulating the group status. In this section, we show in Theorem 1 that when buyers are strategic, a pricing policy without considering buyers’ fairness learning process incurs a linear regret lower bound of $\Omega(T)$. We now present some standard assumptions in the dynamic pricing literature. In later sections, we will show that our proposed pricing policy achieve a sub-linear regret under the same assumptions.
@@
The prices $p_{0t},p_{1t}\in(0,B)$, the feature $\|\bm{x}_{t}\|_{2}\leq x_{max}$, the demand parameters $a_{min}\leq|\alpha_{j}|\leq a_{max}$, $\|\bm{\beta}_{j}\|_{1}\leq b_{max}$ for $j=0,1$, for some positive constants $B,x_{max},a_{min},a_{max},b_{max}$.
Assumption 1 indicates that the price, features and demand parameters are all bounded. The bounded assumptions are practical and also commonly used in pricing literature (Luo et al., 2022, 2024; Zhao et al., 2023; Wang et al., 2024; Fan et al., 2024).
@@
Feature vectors are generated from a fixed distribution. The minimum eigenvalue of the second-moment matrix $\Sigma_{x}=\mathbb{E}(\bm{x}_{t}\bm{x}_{t}^{\top})$ is positive, i.e., $\lambda_{min}(\Sigma_{x})>0.$
Assumption 2 is mild and requires that no features are perfectly collinear in order to identify the true demand parameters (Fan et al., 2024; Liu et al., 2024; Chai et al., 2024; Zhao et al., 2024).
@@
Theorem 1 shows that the fair pricing policy without buyers’ fairness learning incurs a linear regret lower bound of $\Omega(T)$, indicating the importance of studying the buyers’ learning process of fairness when designing fair pricing policies. Intuitively, when buyers from the disadvantage group do not learn the price difference, they always report a false group status. In this case, there exists a problem instance such that it always incurs $\Omega(1)$ regret whenever a buyer from the disadvantage group enters the market. Then, the cumulative regret at time $T$ is at least $\Omega(T)$. Motivated from this, in Section 3, we develop a new fair dynamic pricing policy by taking buyers’ strategic behaviors and buyers’ fairness learning process into consideration.

### v2

@@ Abstract
Contextual pricing strategies are prevalent in online retailing, where the seller adjusts prices based on products’ attributes and buyers’ characteristics. Although such strategies can enhance seller’s profits, they raise concerns about fairness when significant price disparities emerge among specific groups, such as gender or race. These disparities can lead to adverse perceptions of fairness among buyers and may even violate thelaw and regulation. In contrast, price differences can incentivize disadvantaged buyers to strategically manipulate their group identity to obtain a lower price. In this paper, we investigate contextual dynamic pricing with fairness constraints, taking into account buyers’ strategic behaviors when their group status is private and unobservable from the seller. We propose a dynamic pricing policy that simultaneously achieves price fairness and discourages strategic behaviors. Our policy achieves an upper bound of $O(\sqrt{T}+H(T))$ regret over $T$ time horizons, where the term $H(T)$ captures the effect of buyers’ perceived price difference. When buyers are able to learn the fairness of the price policy, this upper bound reduces to $O(\sqrt{T})$. We also prove an $\Omega(\sqrt{T})$ regret lower bound of any pricing policy under our problem setting. We support our findings with extensive experimental evidence, showcasing our policy’s effectiveness. In our real data analysis, we observe the existence of price discrimination against race in the loan application even after accounting for other contextual information. Our proposed pricing policy demonstrates a significant improvement, achieving an average reduction of 30.71% in regret compared to the benchmark policy.
@@ 1 Introduction
Practical scenarios often involve price discrimination against specific groups, even after controlling for contextual information (Bocian et al., 2008; Zhang, 2018; Bartlett et al., 2022; Butler et al., 2023). One such example is the mortgage market. Our motivation comes from the Home Mortgage Disclosure Act (HMDA) data [^4] https://ffiec.cfpb.gov/data-publication/2022, where Popick (2022) found that minority applicants pay higher interest rates than applicants from the majority race, even after controlling for credit risk, underscoring the unfairness of such practices. We refer to Section 6 for more discussion of this HMDA dataset. Another instance is from the auto repair industry, where discrimination against female customers is prevalent (Busse et al., 2017). Nationwide, women are commonly charged more than men for the same auto repair work. In Los Angeles, 20% of auto shops surveyed quoted higher prices for women. On average, women are charged 8% more than men for repair jobs across the country [^5] https://abc7.com/women-overcharged-in-auto-repair-shops-charges-charged-for-repairs-pal/1660671/.
Unfairness in the pricing policy not only can lead to losses for the seller, but also can provoke strategic behaviors among buyers. In personalized pricing, buyers should not be able to easily obtain prices intended for a different consumer group, or if they can, the process should be sufficiently costly (Lukacs et al., 2016). Our study delves into buyers’ strategic behaviors, where such behavior is defined as buyers pretending to belong to an alternative group, incurring a fixed cost in the process (Li and Li, 2023). As revealed in our analysis of HMDA data in Section 6, minority applicants tend to pay higher interest rates compared to the majority group. In response, applicants from the minority group may strategically manipulate their identity to appear as members of the majority group, aiming for a lower interest rate. Consistent with this concern, Dobre and Witzen (2023) showed that measured racial disparities in mortgage outcomes are sensitive to whether race is identified by lenders or by applicants: the estimated gaps increase by 11%-14% for minority borrowers when lender-identified race is used. This discrepancy suggests that misreporting in HMDA is nontrivial and that strategic identity manipulation is empirically relevant. Practical instances of such strategic behaviors exist in reality. For instance, homeowners from the minority race asked friends from the majority race to pretend to be homeowners during property appraisals, leading to a significant increase in property value [^6] https://www.cnn.com/2021/12/09/business/black-homeowners-appraisal-discrimination-lawsuit/index.html
@@
. In cases like these, sellers often obtain buyers’ race information on the basis of buyers’ provided information, visual observation or surname [^7] https://www.consumerfinance.gov/rules-policy/regulations/1003/b/. Disadvantaged buyers may seek the assistance of more advantaged friends to appear in the process of purchasing. Other strategic behaviors include manipulating device information in the case of Orbitz’ s price discrimination against Mac users (Mattila and Choi, 2014), or forging a student ID if there is a discount for students. When significant unfairness arises, buyers may be motivated to manipulate their group membership to gain access to a lower price, even if it incurs additional costs.
@@ 1.1 Our Contribution
To address the aforementioned contextual dynamic pricing problem with fair-minded and strategic buyers, we propose a fairness-aware pricing policy designed to deter buyers’ strategic actions by fostering a favorable fairness perception. We focus on fairness-aware dynamic pricing when buyers exhibit boundedly rational behaviors, and do not aim to characterize a full Nash equilibrium of the dynamic pricing game under uncertainty.
@@ 1.2 Literature Review
Dynamic Pricing with Fairness. Dynamic pricing has been an active research area in operations research and machine learning (Luo et al., 2022; Fan et al., 2024). In the online pricing realm, Xu et al. (2023); Cohen et al. (2025); Chen et al. (2026a) have explored the non-contextual pricing problem with fairness constraint. Specifically, Cohen et al. (2025) and Chen et al. (2026a) examined pricing problems where the identical deterministic price is offered within each buyer group, while Xu et al. (2023) introduced random prices generated by probability distributions within each buyer group. Consequently, in these studies, buyers from the same group are subject to the same price or prices from the same probability distribution. In contrast, our work integrates contextual information into the pricing policy, offering prices based on features while simultaneously ensuring fairness among buyers with the same features from different groups. The work by Chen et al. (2026b) studied the contextual pricing problem with fairness constraint. In all the aforementioned literature, the true group status is observable by the seller while the true group status is unobservable in our paper. Moreover, we consider inequity-averse buyers who actively seek lower prices by manipulating their group status - an aspect not addressed in Xu et al. (2023); Chen et al. (2026b, a); Cohen et al. (2025). Another critical difference is that the aspect of how buyers learn about price fairness is overlooked in Xu et al. (2023); Chen et al. (2026b, a); Cohen et al. (2025). In contrast, our approach delves into the intricacies of how buyers learn about fairness. The existence of buyers’ strategic behaviors is well motivated in many pricing applications, rendering existing pricing tools not applicable. To address this challenge, we must devise new tools capable of handling both strategic behaviors and the learning process of fairness.
Dynamic Pricing with Strategic Buyers. Existing literature on pricing with strategic buyers has primarily focused on timing (Chen and Farias, 2018), untruthful bidding in pricing and auction design (Amin et al., 2014; Mohri and Munoz, 2015), and feature manipulation (Liu et al., 2025). Other literature also exists on feature manipulation within classification (Hardt et al., 2016; Dong et al., 2018; Chen et al., 2020; Ghalme et al., 2021; Bechavod et al., 2021). Our work specifically addresses strategic behaviors related to feature manipulation, closely related to the study by Liu et al. (2025). However, the policy presented by Liu et al. (2025) incurs a social loss as it cannot effectively curb the futile strategic manipulation. In contrast, our paper employs fairness as a tool to discourage strategic behaviors. Furthermore, Liu et al. (2025) could not enforce fairness in pricing policies and hence is not applicable to address price discrimination.
@@ 1.3 Organization
The remainder of the paper is organized as follows. In Section 2, we introduce the new problem and the necessary components: fairness and strategic behaviors. In Section 3, we propose the fairness-aware pricing policy with strategic buyers. In Section 4, we provide the theoretical analysis. In Section 5, we conduct simulation studies to examine our proposed policy. A real data analysis is provided in Section 6, followed by some discussion of future work in Section 7. The supplementary materials include additional information and proofs.
@@ 2 Problem Formulation
We first introduce the setting of the contextual dynamic pricing. At each time $t$, a buyer with feature $\bm{x}_{t}\in\mathbb{R}^{d}$ and a group status $G_{t}\in\{0,1\}$ enters the market. The true group status of each buyer is considered as a private type, which is unobservable by the seller. The buyer can manipulate it to the other group by incurring some cost. We denote $C_{0}>0$ as the fixed unit cost of manipulation, averaged over the total number of products purchased. Upon receiving $\bm{x}_{t}$ and a reported group status $G_{t}^{\prime}$, which may be different from $G_{t}$, the seller offers a price $p_{t}$. At time $t$, the demand of a buyer with feature $\bm{x}_{t}$ and group status $G_{t}=j\in\{0,1\}$ is
$y_{jt}=\alpha_{j}p_{t}+\bm{\beta}_{j}^{\top}\widetilde{\bm{x}}_{t}+\epsilon_{t},$ (1)
where $\alpha_{j}<0$ and $\bm{\beta}_{j}$ are unknown parameters, and $\widetilde{\bm{x}}_{t}=(1,\bm{x}_{t}^{\top})^{\top}$. For convenience, we denote the unknown parameters by $\bm{\theta}_{j}=(\alpha_{j},\bm{\beta}_{j}^{\top})^{\top}$. The demand depends on the parameter $\bm{\theta}_{j}$ corresponding to the true group status rather than those associated with the reported group status. Here, $y_{jt}$ is the demand quantity, such as the amount of the loan the borrower wants to apply for. This linear demand model has been widely considered in the pricing literature (Simchi-Levi and Wang, 2023; Chen et al., 2024). We use $\bm{x}_{t}$ to denote the feature vector (not including group status $G_{t}$) and is unchangeable by buyers. Without loss of generality, $\mathbb{E}\bm{x}_{t}$ is normalized to $\bm{0}$ (Cai et al., 2023). The noise $\epsilon_{t}$ is an independent and identical distributed $(i.i.d.)$ $\sigma^{2}_{\epsilon}$-sub-Gaussian variable and $\mathbb{E}(\epsilon_{t}|\bm{x}_{t},p_{t})=0$.
Now we consider the dynamic pricing problem with price fairness. Let $p_{0t}$ and $p_{1t}$ denote the prices intended for group 0 and group 1, respectively. To incorporate the price fairness, we consider a price constraint, $|p_{0t}-p_{1t}|\leq\delta$, where $\delta\geq 0$ is the parameter for the fairness level that is selected by the seller to meet the internal goal of the company or satisfy regulatory requirements. This price constraint indicates that the price difference of the buyers from two different groups should not exceed $\delta$ after controlling other features.
@@
The objective function of the weighted revenue in (2) is an extension of Xu et al. (2023) from the non-contextual setting to the contextual setting. It also includes Cohen et al. (2025); Chen et al. (2026a) as a special case with $q=1/2$. In the main paper, we focus on the two-group case and provide an extension to the multi-group setting in Appendix S.5.
The constrained optimization problem (2) serves as a full-information benchmark for evaluating our pricing policy. By solving (2) as detailed in Appendix S.9.1, we obtain the optimal prices $p_{0t}^{*}=p_{0}^{*}(\bm{x}_{t})$ and $p_{1t}^{*}=p_{1}^{*}(\bm{x}_{t})$, i.e., for $j=0,1$,
@@
Now, we discuss buyers’ strategic behavior. The group status of the buyer is considered as a private type and is unobserved. The buyers can strategically deceive the seller about the group status to pursue a lower price. Remind that the cost of manipulating group status is $C_{0}>0$, which is public information (Li and Li, 2023). The cost of misreporting group status reflects the efforts and time a buyer undertakes to successfully mimic the other type. Assuming a known switching cost is standard in the strategic learning literature (Shavit et al., 2020; Ghalme et al., 2021; Bechavod et al., 2022; Harris et al., 2022a). While the exact value of $C_{0}$ may not be directly observable in practice, it can be reasonably approximated. For example, if manipulation requires assistance from a member of the advantaged group, publicly available wage data could serve as a practical proxy for estimating the cost. Given that group 0 is the discriminated group, only the buyers from group 0 are likely to manipulate the group status. Let $\widehat{p}_{0}(\bm{x})$ and $\widehat{p}_{1}(\bm{x})$ be the prices for group 0 and group 1 that the buyer has estimated using the history data. For a buyer from group 0 with feature $\bm{x}$, the total estimated cost is $\widehat{p}_{0}(\bm{x})$ without group manipulation and $\widehat{p}_{1}(\bm{x})+C_{0}$ after group manipulation. Therefore, the buyer from group 0 will strategically report the group status as
@@
Remark 1.
The fairness level $\delta$ in (2) is fixed and chosen to be less than $C_{0}$. In practice, buyers can learn a $\widehat{\delta}$ that approximates $\delta$, but this estimate may be either higher or lower than $\delta$. If the seller sets $\delta=C_{0}$, then $\widehat{\delta}$ may exceed $C_{0}$, making misreporting potentially worthwhile. To robustly deter such strategic behaviors in the presence of estimation error, the seller needs to choose $\delta<C_{0}$. While one could treat $\delta$ as a time-varying strategy and achieve $\lim_{t\rightarrow\infty}\delta_{t}=C_{0}$, which could potentially produce higher profits. In that case, a different full-information benchmark is used to evaluate the pricing policy. This would significantly change the scope of the problem and increase the complexity in the analysis. Since our focus is to evaluate regret rather than identifying which full-information benchmark yields the highest profit, we treat $\delta$ as a fixed parameter rather than a strategy of the seller.
@@ 2.1 Linear Regret for Existing Fair Pricing Policy
Consideration of buyers’ fairness learning is critical in the dynamic pricing with strategic buyers. Existing fair pricing policies (Chen et al., 2026b, a; Cohen et al., 2025) do not consider buyers’ strategic behaviors and ignore buyers’ fairness learning process. In this case, even the seller provides prices with fairness constraints, the disadvantaged buyers always act strategically by manipulating the group status. In this section, we show in Theorem 1 that when buyers are strategic, a pricing policy without considering buyers’ fairness learning process incurs a linear regret lower bound of $\Omega(T)$. We now present some standard assumptions in the dynamic pricing literature. In later sections, we will show that our proposed pricing policy achieve a sub-linear regret under the same assumptions.
@@
For each group $j=0,1$, the following conditions hold: The price $p_{jt}\in(0,B)$, the feature vector $\|\bm{x}_{t}\|_{2}\leq x_{max}$, the demand parameters $a_{min}\leq|\alpha_{j}|\leq a_{max}$, $\|\bm{\beta}_{j}\|_{1}\leq b_{max}$, for some positive constants $B,x_{max},a_{min},a_{max},b_{max}$.
Assumption 1 indicates that the price, features and demand parameters are all bounded. The bounded assumptions are practical and also commonly used in pricing literature (Luo et al., 2022, 2024; Zhao et al., 2023; Wang et al., 2024; Fan et al., 2024). In particular, the condition that $|\alpha_{j}|$ is lower bounded by a positive constant is adopted in prior works (Nambiar et al., 2019; Simchi-Levi and Wang, 2023; Chen et al., 2024), as it guarantees that price has a non-negligible influence on demand. This is essential for the effectiveness of pricing optimization. We define the parameter sets $\mathcal{A}=\{\alpha\in\mathbb{R}:a_{min}\leq|\alpha|\leq a_{max},\alpha<0\}$ and $\mathcal{B}=\{\bm{\beta}\in\mathbb{R}^{d}:\|\bm{\beta}\|_{1}\leq b_{max}\}$. Consequently, we assume $\alpha_{j}\in\mathcal{A}$ and $\bm{\beta}_{j}\in\mathcal{B}$ for $j=0,1$.
@@
Feature vectors are generated independently from a fixed distribution. The minimum eigenvalue of the second-moment matrix $\Sigma_{x}=\mathbb{E}(\bm{x}_{t}\bm{x}_{t}^{\top})$ is positive, i.e., $\lambda_{min}(\Sigma_{x})>0.$
Assumption 2 is mild and requires that no features are perfectly collinear in order to identify the true demand parameters (Fan et al., 2024; Chai et al., 2024; Liu et al., 2025; Zhao et al., 2026).
@@
Theorem 1 shows that the fair pricing policy without buyers’ fairness learning incurs a linear regret lower bound of $\Omega(T)$, indicating the importance of studying the buyers’ learning process of fairness when designing fair pricing policies. Intuitively, when buyers from the disadvantage group do not learn the price difference, they always report a false group status. In this case, there exists a problem instance such that it always incurs $\Omega(1)$ regret whenever a buyer from the disadvantage group enters the market. Then, the cumulative regret at time $T$ is at least $\Omega(T)$. To address this, in Section 3, we develop a new fair dynamic pricing policy by taking buyers’ strategic behaviors and buyers’ fairness learning process into consideration.